Calculate a 95% confidence interval for the standard deviation of Total Cost for all customers (assuming that Total Cost follows normal distribution).

n 400

Variance $9,481.51

Std Dev 97.37

Sample mean $153.82
Answer #1

Solution :

Given that,

= $153.82

s = $97.37

n = 400

Degrees of freedom = df = n - 1 = 400 - 1 = 399

At 95% confidence level the t is,

= 1 - 95% = 1 - 0.95 = 0.05

/ 2 = 0.05 / 2 = 0.025

t /2,df = t0.025,399 = 1.966

Margin of error = E = t/2,df * (s /n)

= 1.966* (97.37 / 400)

= 9.571

The 95% confidence interval estimate of the population mean is,

- E < < + E

153.82 - 9.571 < < 153.82 + 9.571

144.249 < < 163.391

(144.249, 163.391)